# Non-linear conductances of Galton–Watson trees and application to the (near) critical random cluster model

### v1

@@ 2.3.2 Near-supercritical connection probability
Once the critical point $\beta_{c}$ is identified, it is also natural to consider the near-critical regime, i.e. take an inverse temperature that may depend on $n$, $\beta=\beta_{n}$, with $\lim_{n\to\infty}\beta_{n}=\beta_{c}$. This is why we allowed in (1.8) the resistance $R_{n}=\psi_{\mathtt{q}}(\beta_{n})$ to also depend on $n$. In fact, Theorem 2.6 can be generalized to the whole near-critical regime. Let us however stress that in (2.5), the function $g_{\beta}=g_{\beta_{n}}$ also depends on $n$, but the bounds (1.5) and the expansion (2.6) are uniform in $n$ (as soon as $\beta_{n}\to\beta_{c}$), so Theorem 2.6 can indeed be extended to the whole near-critical regime as a consequence of Theorems 1.8 and 1.10.
For simplicity of the exposition (and for later use), let us state only the case where $\lim_{n\to\infty}\beta_{n}=\beta_{c}$, but in the near-supercritical regime, that is when $m\psi_{\mathtt{q}}(\beta_{n})-1\gg\frac{1}{n}$. Note that we have,
@@
since $\psi_{\mathtt{q}}^{\prime}(\beta_{c})=\frac{\mathtt{q}\,e^{\beta_{c}}}{(e^{\beta_{c}}+\mathtt{q}-1)^{2}}=\frac{m-1}{m^{2}\mathtt{q}}(m+\mathtt{q}-1)$, recalling also that $e^{\beta_{c}}=\frac{m+\mathtt{q}-1}{m-1}$, see Remark 2.5. Using also (1.10) when $m\psi_{\mathtt{q}}(\beta_{n})-1\gg\frac{1}{n}$ we have
$a_{n}:=\sum_{k=1}^{n}(m\psi_{\mathtt{q}}(\beta_{n}))^{-ks}\sim\frac{1}{(m\psi_{\mathtt{q}}(\beta_{n}))^{s}-1}\sim s^{-1}(m\psi_{\mathtt{q}}(\beta_{n})-1)^{-1}\,,$
we directly obtain the following result from Theorem 1.8-Proposition 1.9 and Theorem 1.10.
@@
Let $(\beta_{n})_{n\geqslant 0}$ be such that $\lim_{n\to\infty}\beta_{n}=\beta_{c}$, with also $\lim_{n\to\infty}n(\beta_{n}-\beta_{c})=+\infty$. Define $\pi_{n}(\beta_{n}):=\bar{\mathbb{P}}_{\mathtt{p}_{n},\mathtt{q}}^{\bar{T}_{n}}(\rho\leftrightarrow\partial T_{n})$, with $\mathtt{p}_{n}=1-e^{-\beta_{n}}$ (which goes to $\mathtt{p}_{c}$). Then, letting $q=2$ if $\mathtt{q}\in(0,2)$ and $q=3$ if $\mathtt{q}=2$, we have the following:
  • If $\mathbf{E}[Z^{q}]<+\infty$, then we have that, in $L^{q}$,
    $\lim_{n\to\infty}\frac{\pi_{n}(\beta_{n})}{(\beta_{n}-\beta_{c})^{\frac{1}{q-1}}}=\tilde{\alpha}_{\mathtt{q}}W\,,\qquad\text{ with }\ \tilde{\alpha}_{\mathtt{q}}:=\begin{cases}\frac{m+\mathtt{q}-1}{\mathtt{q}-2}\mathbf{E}[W^{2}]^{-1}&\text{ if }\mathtt{q}\in(0,2)\,,\\[5.0pt] \sqrt{\frac{3m}{2}}\,\mathbf{E}[W^{3}]^{-1/2}&\text{ if }\mathtt{q}=2\,.\end{cases}$
  • If $c_{1}L(x)x^{q-\alpha}\leqslant\mathbf{E}[(Z\wedge x)^{q}]\leqslant c_{2}L(x)x^{q-\alpha}$ for some $\alpha\in(1,q]$, some slowly varying function $L(\cdot)$ and constants $c_{1},c_{2}>0$, then
@@
Notice that in Theorem 2.8, one can take $\beta_{n}\downarrow\beta_{c}$ arbitrarily slowly. We can therefore deduce the following critical behavior for the limiting survival probability $\pi(\beta)$ of the RCM on a quenched Galton–Watson tree.
@@
Recall the definition (2.7) of $\pi(\beta)=\pi(\beta,T)$. Then, letting $q=2$ if $\mathtt{q}\in(0,2)$ and $q=3$ if $\mathtt{q}=2$, we have the following critical behavior:
  • If $\mathbf{E}[Z^{q}]<+\infty$, then $\pi(\beta)\sim\tilde{\alpha}_{\mathtt{q}}W\,(\beta-\beta_{c})^{\frac{1}{q-1}}$, with $\tilde{\alpha}_{\mathtt{q}}$ as in Theorem 2.8.
  • If $\mathbf{E}[(Z\wedge x)^{q}]\asymp L(x)x^{q-\alpha}$, then $\pi(\beta)\asymp h^{-1}(\beta-\beta_{c})$, with $h^{-1}(\cdot)$ an asymptotic inverse of $L(1/x)x^{\alpha-1}$ as $x\downarrow 0$.
Let us stress that Corollary 2.9 is related to existing results. For instance, [17] gives the critical behavior of the Ising model (with external field) on a tree-like graph. Our techniques are quite different, and we improve the results here by giving a sharp behavior in the case where $\mathbf{E}[Z^{q}]<+\infty$ (with the correct random constant); and treating a more general set-up in the heavy-tail case. Let us also note that [28] treats the survival probability of Bernoulli percolation on a GW tree and our Corollary 2.9 improves their main result in two directions: (i) we recover the first-order asymptotic with a weaker moment condition ($\mathbf{E}[Z^{2}]<+\infty$ instead of $\mathbf{E}[Z^{3+\eta}]<+\infty$); (ii) we treat the case of a GW with heavy tails.
@@ 3.1 Useful $L^{q}$ inequalities for sums of independent random variables
One of the main inequalities that we will use is Lemma 1.4 in [24], which mostly relies on the Marcinkiewicz–Zigmund inequality (see e.g.  [10, Ch. 10.3]).
@@
Now, using Lemma 3.1 conditionnally on $N$, we have that
@@
Using that $\|X-\mathbf{E}[X]\|_{q}\leqslant\|X\|_{q}+\mathbf{E}[X]$, this directly concludes (3.1), since $\sup_{k}\|W_{k}\|_{\alpha}=\|W\|_{\alpha}$ for any $\alpha\in[1,q]$ ($(W_{k}^{\alpha})_{k\geqslant 0}$ is a submartingale), and also $m^{-k(1-\frac{1}{q}\theta_{q})}\leqslant 1$ since $\theta_{q}<q$.
@@ 3.2 Supercritical branching processes with (truncated) heavy tails
Let $\mu$ be the reproduction law of a super-critical branching process, let $Z\sim\mu$ and note $m:=\mathbf{E}[Z]$ its mean. For $t>1$ some (large) fixed parameter, we define $\tilde{Z}:=Z\wedge t$ with distribution denoted $\tilde{\mu}$ and assume that $t$ is large enough so that $\tilde{m}:=\mathbf{E}[\tilde{Z}]>1$. We define a truncated branching process with reproduction law $tilde{\mu}$. As above, let $(Z_{k})_{k\geqslant 0}$, resp. $\tilde{Z}_{k}$, be a branching process with offspring distribution $Z\sim\mu$, resp. $\tilde{Z}\sim\tilde{\mu}$, and denote $W_{k}:=\frac{1}{m^{k}}Z_{k}$ and $\tilde{W}_{k}:=\frac{1}{\tilde{m}^{k}}\tilde{Z}_{k}$ the corresponding martingales.
@@ 4 Estimates on moments of $C_{n}$ in the case $\mathbf{E}[Z^{q}]<+\infty$
Let us denote $C_{n}(u)$ the conductance of the subtree $T_{n}(u)$, equipped with the same resistances. Then setting $\phi(u)=R_{u}C_{n}(u)$ we have the following recursion: for any $u\notin\partial T_{n}$,
@@ 4.1.1 Upper bound on $\mathbf{E}[C_{n}]$
We start from the recursion (4.1). We proceed as in [3, §6.2.1] (see also [9, Lem. 3.1] in the case $p=q=2$). Using the branching property and the fact that $R_{u}=R^{-|u|}$, we get that for any $u\notin\partial T_{n}$ and $v\leftarrow u$,
$\mathbf{E}[\phi(u)]=mR_{n}\,\mathbf{E}\left[\frac{\phi(v)}{(1+\phi(v)^{s})^{1/s}}\right]\leqslant mR_{n}\,\frac{\mathbf{E}[\phi_{p}(v)]}{(1+\mathbf{E}[\phi(v)]^{s})^{1/s}}\,,$
@@ 4.1.2 Lower bound on $C_{n}$, $\mathbf{E}[C_{n}]$
For the lower bound on $C_{n}=\mathcal{C}_{p}(\rho\leftrightarrow\partial T_{n})$, we actually prove an upper bound for the $p$-resistance $\mathcal{R}_{p}(\rho\leftrightarrow\partial T_{n})$; we proceed as in [3, §6.2.2]. Using Thomson’s principle (1.2), we obtain an upper bound on the $p$-resistance simply by computing the energy of a well-chosen unit flow $\theta$ from $\rho$ to $\partial T_{n}$. The uniform flow $\hat{\theta}$ on $T$ is a natural choice (as in [32, Lem. 2.2], see also [9, Lem. 3.3]): define
@@
where $Z_{n}(v)$ is the number of descendants of $v$ at generation $n$ and $Z_{n}:=Z_{n}(\rho)$. Then, $\hat{\theta}(u,v)$ is easily seen to be a unit flow from the root to the leaves of $T_{n}$, and we have
@@
Defining $W_{n}(v):=\frac{Z_{n}(v)}{\mathbf{E}[Z_{n}(v)]}=m^{-(n-k)}Z_{n}(v)$ for $|v|=k$ and $W_{n}:=W_{n}(\rho)$, using that $R_{u}=R^{-|u|}$ and recalling that $s=q-1$, we can rewrite the upper bound as
@@
Note that more generally, the proof shows that, for any $u\in T_{n}$ with $|u|=k$,
@@ 4.2 Control of the higher moments of $C_{n}$
The proof would work similarly as in [9, Lem. 3.2] if $r$ were an integer, but we aim at generalizing their result here (both to non-integer moments $r$, and to a $p$-conductance with $p\neq 2$). We take inspiration from [24, Prop. 1.3], which deals with the rate of convergence in $L^{q}$ for the usual martingale $W_{n}$.
@@
Recall that we define $\phi(u):=R_{u}C_{n}(u)=(R_{n})^{-|u|}C_{n}(u)$. Now, using (4.1), we have the general upper bound
@@
where we have also used that $\varphi_{k_{r}}\leqslant(c_{2})^{k_{r}}\varphi_{0}$, see Remark 4.1. We can then iterate this inequality, applying it to $\|\phi_{k}\|_{r},\|\phi_{2k}\|_{r}$, etc. Letting $n=k_{r}n_{r}+j_{r}$ with $10\leqslant j_{q}\leqslant k_{r}-$ 1, we have
@@ 4.3 Control of the ratios $\varphi_{k}/\varphi_{k-1}$
We start with the iteration (4.1) which defines $\phi(u)$, with $R_{u}=(R_{n})^{-|u|}$ Notice that we can write the iteration as
@@
Taking the expectation, we therefore get that if $|u|=k-1$,
@@ 5 Convergence of the normalized conductance
The method we use is analogous to that in [9, §5], with some modifications to obtain the convergence in $L^{q}$ with only a finite $q$-moment assumption. We start by proving the $L^{1}$ convergence; then we upgrade it to a $L^{q}$ convergence. We then conclude the section by proving the almost sure convergence in the case $R_{n}\equiv R$. We conclude the section by proving the sharp asymptotic of Proposition 1.9.
Preliminary observations and notation.

### v2

@@ 2.3.2 Near-supercritical connection probability
Once the critical point $\beta_{c}$ is identified, it is also natural to consider the near-critical regime, i.e. take an inverse temperature that may depend on $n$, $\beta=\beta_{n}$, with $\lim_{n\to\infty}\beta_{n}=\beta_{c}$. This is why we allowed the resistances $R_{n}=\psi_{\mathtt{q}}(\beta_{n})$ in (1.8) to also depend on $n$. In fact, Theorem 2.6 can be generalized to the whole near-critical regime. Let us however stress that in (2.5), the function $g_{\beta}=g_{\beta_{n}}$ also depends on $n$, but the bounds (1.9) and the expansion (2.6) are uniform in $n$ (as soon as $\beta_{n}\to\beta_{c}$), so Theorem 2.6 can indeed be extended to the whole near-critical regime as a consequence of Theorems 1.8 and 1.10.
For simplicity of the exposition (and for later use), let us state only the case where $\lim_{n\to\infty}\beta_{n}=\beta_{c}$, but in the near-supercritical regime, that is when $m\psi_{\mathtt{q}}(\beta_{n})-1\gg\frac{1}{n}$. Note that we have
@@
since $\psi_{\mathtt{q}}^{\prime}(\beta_{c})=\frac{\mathtt{q}\,e^{\beta_{c}}}{(e^{\beta_{c}}+\mathtt{q}-1)^{2}}=\frac{m-1}{m^{2}\mathtt{q}}(m+\mathtt{q}-1)$, recalling also that $e^{\beta_{c}}=\frac{m+\mathtt{q}-1}{m-1}$, see Remark 2.5. Using that when $m\psi_{\mathtt{q}}(\beta_{n})-1\gg\frac{1}{n}$ we have
$a_{n}:=\sum_{k=1}^{n}(m\psi_{\mathtt{q}}(\beta_{n}))^{-ks}\sim\frac{1}{1-(m\psi_{\mathtt{q}}(\beta_{n}))^{-s}}\sim s^{-1}(m\psi_{\mathtt{q}}(\beta_{n})-1)^{-1}\,$,
we get
$a_{n}\sim\frac{m\mathtt{q}}{s(m-1)(m+\mathtt{q}-1)}\,(\beta_{n}-\beta_{c})^{-1}\,,\quad\text{ as }\beta_{n}\to\beta_{c}\,.$
We therefore directly obtain the following result from Theorem 1.8-Proposition 1.9 and Theorem 1.10.
@@
Let $(\beta_{n})_{n\geqslant 0}$ be such that $\lim_{n\to\infty}\beta_{n}=\beta_{c}$, with also $\lim_{n\to\infty}n(\beta_{n}-\beta_{c})=+\infty$. Define $\pi_{n}(\beta_{n}):=\bar{\mathbb{P}}_{\mathtt{p}_{n},\mathtt{q}}^{\bar{T}_{n}}(\rho\leftrightarrow\partial T_{n})$, with $\mathtt{p}_{n}=1-e^{-\beta_{n}}$ (which goes to $\mathtt{p}_{c}$). Then, letting $s=1$ if $\mathtt{q}\in(0,2)$ and $s=2$ if $\mathtt{q}=2$, we have the following:
  • If $\mathbf{E}[Z^{1+s}]<+\infty$, then we have that almost surely and in $L^{1+s}$,
    $\lim_{n\to\infty}\frac{\pi_{n}(\beta_{n})}{(\beta_{n}-\beta_{c})^{1/s}}=\tilde{\alpha}_{\mathtt{q}}W\,,\qquad\text{ with }\ \tilde{\alpha}_{\mathtt{q}}:=\begin{cases}\frac{2(m+\mathtt{q}-1)}{\mathtt{q}(2-\mathtt{q})}\mathbf{E}[W^{2}]^{-1}&\text{ if }\mathtt{q}\in(0,2)\,,\\[5.0pt] \sqrt{\frac{3m}{2}}\,\mathbf{E}[W^{3}]^{-1/2}&\text{ if }\mathtt{q}=2\,.\end{cases}$
  • If $c_{1}L(x)x^{1+s-\alpha}\leqslant\mathbf{E}[(Z\wedge x)^{1+s}]\leqslant c_{2}L(x)x^{1+s-\alpha}$ for some $\alpha\in(1,1+s]$, some slowly varying function $L(\cdot)$ and constants $c_{1},c_{2}>0$, then
@@
Notice that in Theorem 2.8, one can take $\beta_{n}\downarrow\beta_{c}$ arbitrarily slowly. We can therefore deduce the following critical behavior for the limiting survival probability $\pi(\beta)$ of the random cluster model on a quenched Galton–Watson tree.
@@
Recall the definition (2.7) of $\pi(\beta)=\pi(\beta,T)$. Then, letting $s=1$ if $\mathtt{q}\in(0,2)$ and $s=2$ if $\mathtt{q}=2$, we have the following critical behavior:
  • If $\mathbf{E}[Z^{1+s}]<+\infty$, then $\pi(\beta)\sim\tilde{\alpha}_{\mathtt{q}}W\,(\beta-\beta_{c})^{1/s}$, with $\tilde{\alpha}_{\mathtt{q}}$ as in Theorem 2.8.
  • If $\mathbf{E}[(Z\wedge x)^{1+s}]\asymp L(x)x^{1+s-\alpha}$, then $\pi(\beta)\asymp h^{-1}(\beta-\beta_{c})$, with $h^{-1}(\cdot)$ an asymptotic inverse of $L(1/x)x^{\alpha-1}$ as $x\downarrow 0$.
Let us stress that Corollary 2.9 is related to existing results. For instance, [19] gives the critical behavior of the Ising model (with external field) on a tree-like graph. Our techniques are quite different, and we improve the results here by giving a sharp behavior in the case where $\mathbf{E}[Z^{1+s}]<+\infty$ (with the correct random constant); and treating a more general set-up in the heavy-tail case. Let us also note that [31] treats the survival probability of Bernoulli percolation on a Galton Watson tree and our Corollary 2.9 improves their main result in two directions: (i) we recover the first-order asymptotic with a weaker moment condition ($\mathbf{E}[Z^{2}]<+\infty$ instead of $\mathbf{E}[Z^{3+\eta}]<+\infty$); (ii) we treat the case of a GW with heavy tails.
@@ 3.1 Useful $L^{q}$ inequalities for sums of independent random variables
One of the main inequalities that we will use is Lemma 1.4 in [27], which mostly relies on the Marcinkiewicz–Zigmund inequality (see e.g. [11, Ch. 10.3]).
@@
Now, using Lemma 3.1 conditionally on $N$, we have that
@@
Using that $\|X-\mathbf{E}[X]\|_{q}\leqslant\|X\|_{q}+\mathbf{E}[X]$, this directly concludes (3.1), since $\sup_{k}\|W_{k}\|_{\alpha}=\|W\|_{\alpha}$ for any $\alpha\in[1,q]$ ($(W_{k}^{\alpha})_{k\geqslant 0}$ is a submartingale), and also $m^{-k(1-\frac{1}{q}\theta_{q})}\leqslant 1$ since we have $\theta_{q}<q$.
@@ 3.2 Supercritical branching processes with (truncated) heavy tails
Let $\mu$ be the reproduction law of a super-critical branching process, let $Z\sim\mu$ and note $m:=\mathbf{E}[Z]$ its mean. For $t>1$ some (large) fixed parameter, we define $\tilde{Z}:=Z\wedge t$ with distribution denoted $\tilde{\mu}$ and assume that $t$ is large enough so that $\tilde{m}:=\mathbf{E}[\tilde{Z}]>1$. We define a truncated branching process with reproduction law $\tilde{\mu}$. As above, let $(Z_{k})_{k\geqslant 0}$, resp. $\tilde{Z}_{k}$, be a branching process with offspring distribution $Z\sim\mu$, resp. $\tilde{Z}\sim\tilde{\mu}$, and denote $W_{k}:=\frac{1}{m^{k}}Z_{k}$ and $\tilde{W}_{k}:=\frac{1}{\tilde{m}^{k}}\tilde{Z}_{k}$ the corresponding martingales.
@@ 4 Estimates on moments of $C_{n}$ in the case $\mathbf{E}[Z^{q}]<+\infty$
For $u\notin\partial T_{n}$, define $\phi(u)=R_{u}\mathcal{C}_{p}(u\leftrightarrow\partial T_{n}(u))$, where we remind that $\mathcal{C}_{p}(u\leftrightarrow\partial T_{n}(u))$ is the effective $p$-conductance on the subtree $T_{n}(u)$, equipped with the original resistances $(R_{u})_{u\in T_{n}}$. We then have the following recursion:
@@ 4.1.1 Upper bound on $\mathbf{E}[C_{n}]$
We start from the recursion (4.1). We proceed as in [3, §6.2.1] (see also [10, Lem. 3.1] in the case $p=q=2$). Using the branching property and the fact that $R_{u}=R^{-|u|}$, we get that for any $u\notin\partial T_{n}$ and $v\leftarrow u$,
$\mathbf{E}[\phi(u)]=mR_{n}\,\mathbf{E}\left[\frac{\phi(v)}{(1+\phi(v)^{s})^{1/s}}\right]\leqslant mR_{n}\,\frac{\mathbf{E}[\phi(v)]}{(1+\mathbf{E}[\phi(v)]^{s})^{1/s}}\,$,
@@ 4.1.2 Lower bound on $C_{n}$, $\mathbf{E}[C_{n}]$
For the lower bound on $C_{n}=\mathcal{C}_{p}(\rho\leftrightarrow\partial T_{n})$, we actually prove an upper bound for the $p$-resistance $\mathcal{R}_{p}(\rho\leftrightarrow\partial T_{n})$; we proceed as in [3, §6.2.2]. Using Thomson’s principle (1.2), we obtain an upper bound on the $p$-resistance simply by computing the energy of a well-chosen unit flow $\theta$ from $\rho$ to $\partial T_{n}$. The uniform flow $\hat{\theta}$ on $T$ is a natural choice (as in [35, Lem. 2.2], see also [10, Lem. 3.3]):
@@
where $Z_{n}(v)$ is the number of descendants of $v$ at generation $n$ and $Z_{n}:=Z_{n}(\rho)$. Then, with this unit flow $\hat{\theta}(u,v)$ from the root to the leaves of $T_{n}$, we have
@@
Defining $W_{n}(v):=\frac{Z_{n}(v)}{\mathbf{E}[Z_{n}(v)]}=m^{-(n-k)}Z_{n}(v)$ for $|v|=k$ and $W_{n}:=W_{n}(\rho)$, using that $R_{u}=R^{-|u|}$ and recalling that s=q1s=q-1, we can rewrite the upper bound as
@@
More generally, the proof shows that, for any $u\in T_{n}$ with $|u|=k$,
@@ 4.2 Control of the higher moments of $C_{n}$
The proof would work similarly as in [10, Lem. 3.2] if $r$ were an integer, but we aim at generalizing their result here (both to non-integer moments $r$, and to a $p$-conductance with $p\neq 2$). We take inspiration from [27, Prop. 1.3], which deals with the rate of convergence in $L^{q}$ for the usual martingale $W_{n}$.
@@
Recall that we define $\phi(u):=R_{u}\mathcal{C}_{p}(u\leftrightarrow\partial T_{n}(u))$, with $R_{u}=(R_{n})^{-|u|}$. Now, using (4.1), we have the general upper bound
@@
where we have also used that $\varphi_{k_{r}}\leqslant(c_{2})^{k_{r}}\varphi_{0}$, see Remark 4.1. We can then iterate this inequality, applying it to $\|\phi_{k}\|_{r},\|\phi_{2k}\|_{r}$, etc. Letting $n=k_{r}n_{r}+j_{r}$ with $0\leqslant j_{q}\leqslant k_{r}-1$, we have
@@ 4.3 Control of the ratios $\varphi_{k}/\varphi_{k-1}$
We start with the iteration (4.1) which defines $\phi(u)$, with $R_{u}=(R_{n})^{-|u|}$. Notice that we can write the iteration as
@@
Taking the expectation, we therefore get that if |u|=k1|u|=k-1,
@@ 5 Convergence of the normalized conductance
The method we use is analogous to that in [10, §5], with some modifications to obtain the convergence in $L^{q}$ with only a finite $q$-moment assumption. We start by proving the $L^{1}$ convergence; then we upgrade it to a $L^{q}$ convergence. We then conclude the section by proving the almost sure convergence in the case $R_{n}\equiv R$ and the sharp asymptotic of Proposition 1.9.
Preliminary observations and notation